# Real-time Localization and Mapping in Architectural Plans with Deviations

### v1

@@ Abstract
Having prior knowledge of an environment boosts the localization and mapping accuracy of robots. Several approaches in the literature have utilized architectural plans in this regard. However, almost all of them overlook the deviations between actual “as-built” environments and “as-planned” architectural designs, introducing bias in the estimations. To address this issue, we present a novel localization and mapping method denoted as deviations-informed Situational Graphs or “diS-Graphs” that integrates prior knowledge from architectural plans even in the presence of deviations. It is based on Situational Graphs (S-Graphs) that merge geometric models of the environment with 3D scene graphs into a multi-layered jointly optimizable factor graph. Our diS-Graph extracts information from architectural plans by first modeling them as a hierarchical factor graph, which we will call an Architectural Graph (A-Graph). While the robot explores the real environment, it estimates an S-Graph from its onboard sensors. We then use a novel matching algorithm to register the A-Graph and S-Graph in the same reference, and merge both of them with an explicit model of deviations. Finally, an alternating graph optimization strategy allows simultaneous global localization and mapping, as well as deviation estimation between both the A-Graph and the S-Graph. We perform several experiments in simulated and real datasets in the presence of deviations. On average, our diS-Graphs outperforms the baselines by a margin of approximately in simulated environments and by in real environments, while being able to estimate deviations up to cm and .
Paper Video: https://www.youtube.com/watch?v=bgPm-sSXZ9g
@@ I Introduction
Prior information from architectural plans can enhance the localization and mapping accuracy of mobile robots. Traditional techniques generally leverage only the metric information available in architectural plans, reducing its robustness in challenging environments. Recent approaches such as 3D scene graphs [1, 2] or Situational Graphs (S-Graphs) [3, 4], represent a robot’s environment in a compact and hierarchical manner, encoding high-level semantic abstractions (for example, walls and rooms) and their relationships (e.g., a set of walls forms a room). Herein, S-Graphs extend 3D scene graphs by merging geometric models of the environment generated by Simultaneous Localization and Mapping (SLAM) approaches with 3D scene graphs into a multi-layered jointly optimizable factor graph. This representation, combined with the prior information extracted from architectural plans, can be used to provide fast and efficient localization.
iS-Graphs [5] follows this direction and shows accurate localization over hierarchical factor graphs using prior information from architectural plans. iS-Graphs extracts elements such as wall-surfaces (planes), walls (two opposite wall-surfaces), rooms, doors, and floors from architectural plans to also model them as a hierarchical factor graph that we call an “Architectural Graph” (A-Graph) [5]. As a robot equipped with a LiDAR navigates the environment, it can detect features of the scene such as walls, rooms, and floors and model them online as an S-Graph. iS-Graphs then performs a hierarchical graph matching and merging to localize the robot within the A-Graph. However, its success is based on the assumption that there are no deviations between the S-Graph (“as-built”) and the A-Graph (“as-planned”). In reality, this is never the case, and the building elements exhibit certain deviations with respect to their planned geometries.
To address this issue, we present in this paper diS-Graphs (Fig. 1), a novel method capable of fusing “as-planned” and “as-built” data even in the presence of deviations. Given an A-Graph of a building, and while an S-Graph is being estimated online from a robot’s sensor readings, our approach simultaneously localizes the robot, maps the environment, and detects and estimates the deviations between the elements of “as-planned” and “as-built” environments. Therefore, our main contributions in this paper are:
A novel real-time localization and mapping algorithm based on Situational Graphs, integrating prior knowledge from architectural plans in the presence of deviations.
A two-stage matching algorithm based on graph structure, identifying deviations between S-Graphs and A-Graphs.
A novel deviation factor between the elements of the S-Graph and A-Graph, for simultaneously estimating global localization and deviations.
@@ II-A Localization and Mapping with Precise Priors
Most localization and mapping techniques using prior information from architectural plans assume that the environments are built precisely according to the plans. One of the most commonly used localization techniques in 2D metric prior maps is Monte Carlo Localization (MCL) [6], [7] but it is not scaleable to large-scale complex environments. Boniardi et al [8] use a technique that scales to more complex environments by aligning a scan-based map with CAD-based floor plans. OGM2PGM [9] also scales to larger environments by converting the 2D floor plan to an occupancy grid map (OGM) and using a pose-graph map (PGM) to localize the robot. UKFL [10] further enhances the localization accuracy using an unscented Kalman filter to localize the robot in 3D metric meshes. Recent techniques such as [11] exploit neural networks to localize the robot using an implicit neural representation of the floor plans. All of the above mentioned techniques primarily rely on geometric information, not using any possible semantic information available in the architectural plans, limiting their ability to reason about the environment beyond geometric features. In addition, inaccuracies or outdated information in the floor plan can significantly affect the performance of these methods.
@@
Semantic-based localization techniques, such as Mendez et al. [12] use semantic cues from architectural plans and sensor information to improve localization accuracy. Boniardi et al. [13] exploit the semantics of the room in architectural plans to do robot localization by matching the detected rooms from sensor data. Wang et al. [14] leverage prelabeled architectural features, such as wall intersections and corners, as landmarks in floor plans, and match them with detection from sensor data to jointly perform mapping and localization. Zimmerman et al. [15], [16] use high-level semantic information in floor plans, derived from object detection, along with geometric data from 2D LiDAR to perform long-term robot localization in floor plans. Huan et al. [17] convert architectural plans into semantically enriched point cloud maps, followed by a coarse-to-fine localization process using ICP. Gao et al. [18] used neural networks to detect vertical elements from floor plans to do LiDAR based localization. These methods are prone to inaccuracies due to misidentification and errors in the pose estimate of semantic elements. Moreover, they do not consider the topological relationship between different semantic elements for a more high-level understanding. Shaheer et al. [5] exploit the topological relationship between semantic elements to localize the robot with respect to architectural plans. However, all the above mentioned approaches assume no deviations between the architectural plans and the actual environment.
II-B Localization and Mapping with Imprecise Priors
@@
Some recent works leverage imprecise floor plans for localization and mapping. Boniardi et al. [19] integrate mapping and localization techniques to take advantage of the information embedded in the CAD drawing, and the real-world observations acquired during navigation, which may not be reflected in the floor plan. Li et al. [20] presented a 2D LiDAR-based localization system in imprecise floor plans using stochastic gradient descent (SGD) with a scan matching algorithm. Chan et al. [21] presented a 2D LiDAR-based localization in floor plans that integrates SLAM with MCL. Blum et al. [22] use neural networks for feature segmentation and combine them with LiDAR data for localization in imprecise floor plans. Although these works can localize the robot in inaccurate floor plans, to the best of our knowledge, none of the existing works can localize the robot while simultaneously providing element-wise deviations between the “as-planned” and the “as-built” environments.
@@ III System Architecture
The overall system architecture of our approach is shown in Fig. 2. The algorithm has two stages: Stage-A and Stage-B, each consisting of multiple processes. In Stage-A, the A-Graph created from the architectural plan is first matched and merged with the S-Graph created by a robot, and then optimized. Stage-A is executed only once to get the initial estimates of the transformation between the two graphs, and potential deviations between its elements. Stage-B is executed periodically, as the new semantic entities are detected by the robot, to match and merge with the A-Graph, until the robot finishes exploring the environment.
III-A Graph Structures
@@
Deviations Informed-Situational Graph (diS-Graph). The result of merging both graphs, estimating the transformation between them and accounting for potential deviations is what we call diS-Graph.
The layers of the graphs are depicted in Fig. 3 and can be detailed as follows:
@@
Keyframes Layer. Only present in S-Graphs, this layer contains as nodes the robot poses in a global reference frame .
@@
Walls Layer. In an A-Graph, this layer’s nodes encode two semantic entities, namely wall-surfaces and walls in the A-Graph global reference . We assume that each wall has two planar wall-surfaces with opposite orientations and the separation between them is equal to the width of the wall. In S-Graphs, this layer contains only the wall-surfaces extracted from 3D LiDAR scans. The keyframes that observe such wall-surfaces are linked to them through pose-plane constraints.
@@
Rooms Layer. In an A-Graph, this layer also encodes two semantic entities, namely Rooms consisting of four wall-surfaces and Doorways . Two rooms constrain a doorway, and a room is constrained by four walls. In S-Graphs, this layer contains rooms comprising either four wall-surfaces or two wall-surfaces, and does not contain doorways.
@@
Floors Layer. In both A-Graph and S-Graph this layer consists of a floor center node represented as , constraining all rooms present at that particular floor level. More details on the type of constraints between the different elements of the graphs can be found in [5].
III-B Graph Matching (Section IV)
@@
The Global Graph Matching in Stage-A provides a unique match, when it exists, between the S-Graph and A-Graph at room and wall-surface levels, accounting for potential deviations. In Stage-B, the Local Graph Matching extends the previously matched elements in the diS-Graph with newly detected elements following an incremental approach.
III-C Graph Merging (Section. V)
@@
Graph merging is performed for the candidates from the matching method from previous subsection. In Stage-A, graph merging involves the registration of the two graphs along with the semantic merging (wall-surfaces and rooms), with and explicit mapping of the deviation factors. Stage-B only performs the semantic merging that incorporates the newly observed entities with explicit deviation factors.
III-D Graph Optimization (Section. VI)
@@
Graph optimization in Stage-A involves two steps of alternating optimization and combined optimization, while Stage-B involves only combined optimization (see Fig. 2). Alternating optimization estimates both the initial transformation between the origins of the S-Graph and the A-Graph, and possible deviations between the matched graph entities. Combined optimization performs a complete graph optimization given the initial global transformation and deviation estimates.
@@ IV Graph Matching
Our graph matching extends the method presented in [5], summarized below for clarity.
Background. In [5], a top-down potential candidate search between the A-Graph and S-Graph is performed by leveraging their hierarchical structures. To assess the overall consistency of each generated candidate, two verification steps are applied iteratively. First, the consistency of the node type and graph structure is verified. Second, geometric consistency (i.e. norm) is maintained over a certain consistency threshold [23]. The hierarchy of the structures is exploited through intra-level (i.e., same level) and inter-level (i.e. different levels) candidate consistency checks in the following steps (S) shown in Fig. 4(b). In S1, consistent room-to-room candidates are generated. In S2, and always consistent with its corresponding room pair, wall-surface-to-wall-surface candidates are generated. In S3, wall-surface candidates are merged with their room-level candidates. In S4, the overall geometrical consistencies of all the remaining candidates are compared. In the case where the score of the best candidate is higher than the second by a certain threshold, it is selected as the final unique match.
@@ IV-A Deviation-based Candidates Search
Non-Deviated Elements Search (Fig. 4(a) first row). To handle potential deviations, our method relaxes the matching criteria by decreasing the consistency thresholds at each level. First, we apply relaxed consistency thresholds for the generation of room-to-room and room-to-wall-surface match candidates, to account for the induced room-center inconsistencies. Then, to exclude deviated wall-surfaces, we increase the threshold for wall-surface-to-wall-surface candidates.
Deviated Elements Detection (Fig. 4(a) second row). To identify the deviated wall-surfaces which were not matched in the first stage but are connected to already matched rooms, we decrease the consistency threshold at wall-surface-to-wall-surface level.
@@
Finally, the geometric consistency score provides a quantification of the probability of the deviation for each room and wall-surface, which is further used in the Graph Merging step (Section V).
@@ IV-B Incremental Matching
To enhance the efficiency of the Graph Matching algorithm in [5], we propose an incremental approach with two stages (associated with stages A and B of the systems architecture of Fig. 2), namely Global Graph Matching (Fig. 4(a) left) and Local Graph Matching (Fig. 4(a) right), each executing the two previously described deviation-based candidate search stages. Until a first unique match has been found, the Global Graph Matching is executed for every new observation in the S-Graph. Afterward, the Local Graph Matching is executed every time the diS-Graph is updated with newly observed rooms and wall surfaces. Here, already-matched elements are excluded from candidate generation, and each assessment of intra-level consistency considers the previously matched elements at the corresponding level.
V Graph Merging
Origin Merging. We first merge the origins of two graphs by introducing a transformation factor . The cost function is defined as:
@@ V Graph Merging
Semantic Merging. Next, we associate the wall-surfaces and rooms of the A-Graph and the S-Graph. To account for and estimate deviations between the two graphs, we introduce deviation factors in graph merging as follows:
Room Merging: To estimate the deviation in the pose of an associated room between the two graphs, we define a deviation factor between the two rooms as , where the cost function is defined as:
@@
Wall-Surface Merging: After associating the rooms of the A-Graph and the S-Graph, we then associate the wall-surfaces of the associated rooms. We define the deviation factor between two wall surfaces as . The cost function to estimate the deviation value is defined as:

### v2

# Tightly Coupled SLAM with Imprecise Architectural Plans

@@ Abstract
Robots navigating indoor environments often have access to architectural plans, which can serve as prior knowledge to enhance their localization and mapping capabilities. While some SLAM algorithms leverage these plans for global localization in real-world environments, they typically overlook a critical challenge: the “as-planned” architectural designs frequently deviate from the “as-built” real-world environments. To address this gap, we present a novel algorithm that tightly couples LIDAR-based simultaneous localization and mapping with architectural plans under the presence of deviations. Our method utilizes a multi-layered semantic representation to not only localize the robot, but also to estimate global alignment and structural deviations between “as-planned” and “as-built” environments in real-time. To validate our approach, we performed experiments in simulated and real datasets demonstrating robustness to structural deviations up to cm and . On average, our method achieves less localization error than baselines in simulated environments, while in real environments, the “as-built” 3D maps show lower average alignment error.
Paper Video: https://www.youtube.com/watch?v=eEZPkcpjWlM
@@ I Introduction
Prior information from architectural plans can be used to enhance the localization and mapping accuracy of mobile robots. However, real-world buildings rarely match their plans perfectly due to construction tolerances and modifications. Incorporating such imprecise prior information into robot navigation pipelines can introduce systematic errors, potentially damaging the localization and mapping accuracy, instead of helping to improve it. To address this issue, it is critical to tightly couple the Simultaneous Localization and Mapping (SLAM) pipeline with the architectural plans. And for this, one needs a unified representation for both “as-planned” and “as-built” environments, that captures not only the geometry but also the architectural semantics. This is what will allow us to match structural elements between the “as-planned” and “as-built” instances and estimate the deviations between the two.
Although there are various environment representation techniques such as occupancy grids [1], surface maps [2], octomap [3] etc, none of them explicitly models the semantic and hierarchical information of environment, which is needed to identify the deviated structural elements. Recent approaches such as 3D scene graphs [4, 5] or Situational Graphs (S-Graphs) [6, 7], represent a robot’s environment in a compact and hierarchical manner, encoding high-level semantic abstractions (for example, walls and rooms) and their relationships (e.g., a set of walls forms a room). Herein, S-Graphs extend 3D scene graphs by merging geometric models of the environment generated by SLAM approaches with 3D scene graphs into a multi-layered jointly optimizable factor graph. This representation, combined with the prior information extracted from architectural plans, can be used to provide fast and efficient localization.
Informed S-Graphs (iS-Graphs) [8] further extend S-Graphs by using architectural plans to provide accurate localization over the resulting hierarchical factor graphs. However, its success is based on the assumption that there are no deviations between the “as-built” and the “as-planned”. In reality, this is never the case, and the building elements exhibit certain deviations with respect to their planned geometries.
The main contribution of this paper is a novel method capable of coupling architectural plans (“as-planned”) and SLAM (“as-built”) data even in the presence of deviations, as shown in Fig. 1. We call this algorithm Deviations Informed Situational Graphs or diS-Graphs in short.
In summary, diS-Graphs performs three interconnected tasks simultaneously:
Coupling of SLAM factor graph with architectural plans to match structural elements (walls, rooms) between planned and built environments.
Globally localizing the robot in imprecise architectural plans.
Detecting and estimating the deviations between the matched structural elements of “as-planned” and “as-built” environments in real-time.
@@ II Related Works
Most localization techniques using prior information from architectural plans assume that the environments are built precisely according to the plans. One of the most commonly used localization techniques in 2D metric prior maps is Monte Carlo Localization (MCL) [9], [10] but it is not scalable to large-scale complex environments. Boniardi et al [11] use a technique that scales to more complex environments by aligning a scan-based map with CAD-based floor plans. OGM2PGM [12] also scales to larger environments by converting the 2D floor plan to an occupancy grid map (OGM) and using a pose-graph map (PGM) to localize the robot. UKFL [13] further enhances the localization accuracy using an unscented Kalman filter to localize the robot in 3D metric meshes. Recent techniques such as [14] exploit neural networks to localize the robot using an implicit neural representation of the floor plans. All of the above mentioned techniques primarily rely on geometric information, not using any possible semantic information available in the architectural plans, limiting their ability to reason about the environment beyond geometric features. This geometric-only approach fails to leverage crucial semantic cues like room types, door locations, and functional spaces that could help disambiguate similar-looking areas and improve localization accuracy. In addition, inaccuracies or outdated information in the floor plan can significantly affect the performance of these methods.
@@
To address these limitations, semantic-based localization techniques, such as Mendez et al. [15] use semantic cues from architectural plans and sensor information to improve localization accuracy. Boniardi et al. [16] exploit the semantics of the room in architectural plans to do robot localization by matching the detected rooms from sensor data. Wang et al. [17] leverage prelabeled architectural features, such as wall intersections and corners, as landmarks in floor plans, and match them with detection from sensor data to jointly perform mapping and localization. Zimmerman et al. [18], [19] use high-level semantic information in floor plans, derived from object detection, along with geometric data from 2D LiDAR to perform long-term robot localization in floor plans. Huan et al. [20] convert architectural plans into semantically enriched point cloud maps, followed by a coarse-to-fine localization process using ICP. Gao et al. [21] used neural networks to detect vertical elements from floor plans to do LiDAR based localization. These methods are prone to inaccuracies due to misidentification and errors in the pose estimate of semantic elements. Moreover, they treat semantic elements in isolation, failing to leverage the rich contextual information embedded in their spatial and functional relationships.
Recent work such as Shaheer et al. [8] exploit the topological relationship between semantic elements to localize the robot with respect to architectural plans. However, all the above mentioned approaches assume no deviations between the architectural plans and the actual environment.
@@
Some recent works leverage imprecise floor plans for localization. Boniardi et al. [22] integrate localization techniques to take advantage of the information embedded in the CAD drawing, and the real-world observations acquired during navigation, which may not be reflected in the floor plan. Li et al. [23] presented a 2D LiDAR-based localization system in imprecise floor plans using stochastic gradient descent (SGD) with a scan matching algorithm. Chan et al. [24] presented a 2D LiDAR-based localization in floor plans that integrates SLAM with MCL. Blum et al. [25] use neural networks for feature segmentation and combine them with LiDAR data for localization in imprecise floor plans. While these works demonstrate robot localization in inaccurate floor plans, they lack the capability to identify element-wise deviations between ”as-planned” and ”as-built” environments.
Despite advances in robot localization using imprecise floor plans, no existing approach combines localization with the estimation of structural deviations between “as-planned” and “as-built” environments, limiting the ability to assess construction accuracy in real time. In this work, we address this limitation by presenting a unified framework that enables both global localization and deviation estimation.
@@ III System Architecture
We propose an algorithm that tightly couples SLAM and architectural plans and jointly optimizes them.
III-A Factor Graph Modelling
Our algorithm models both SLAM and architectural plans as hierarchical factor graphs (see Fig. 3) composed of the following elements:
@@ III-A Factor Graph Modelling
Deviations Informed Situational Graph (diS-Graph). diS-Graph is the result of couppling both A-Graph and S-Graph.
These graphs are organized in the following layers:
Origins Layer. Both S-Graph and A-Graph are defined in their own reference frame. The reference frame of S-Graph and A-Graph are called and respectively. Both the origin frames are coupled via a Transformation factor called .
@@
Keyframes Layer. Only present in S-Graphs, this layer contains as nodes the robot poses .
@@
Walls Layer. In an A-Graph, this layer’s nodes encode two semantic entities, namely wall-surfaces and walls . Every wall-surface is defined by four coefficients representing its normal orientation and distance to origin [6]. We assume that each wall has two planar wall-surfaces with opposite orientations and the separation between them is equal to the width of the wall. In S-Graphs, this layer contains only the wall-surfaces extracted from 3D LiDAR scans. The keyframes that observe such wall-surfaces are linked to them through pose-plane constraints. This layer of both graphs is coupled via wall deviation factors in diS-Graph.
@@
Rooms Layer. In an A-Graph, this layer also encodes two semantic entities, namely Rooms consisting of four wall-surfaces and Doorways . Two rooms constrain a doorway, and a room is constrained by four walls. In S-Graphs, this layer contains rooms comprising either four wall-surfaces or two wall-surfaces, and does not contain doorways. This layer of both graphs is coupled via room deviation factors in diS-Graph.
@@
Floors Layer. In both A-Graph and S-Graph this layer consists of a floor center node represented as , constraining all rooms present at that particular floor level.
More details on the type of constraints between the different elements of the graphs can be found in [8].
III-B The Algorithm
Our algorithm has two stages, as shown in Fig. 2. In Stage-A, the A-Graph is first matched and coupled with the S-Graph, and then jointly optimized. Stage-A runs until it finds the initial match between the two graphs, and provides the initial estimates of both the transformation between graphs and potential deviations between their elements. Afterwards, in Stage-B the joint graph is continuously refined through incremental matching and optimization as the robot navigates the environment, incorporating newly detected semantic entities. This tight coupling enables continuous refinement of both the robot’s localization and the detected structural deviations during navigation.
Both stages of our algorithm consist of multiple processes, described briefly below and detailed in their respective sections:
@@ III-B The Algorithm
Graph Matching (Section IV). The Global Graph Matching in Stage-A provides the first unique match, when it exists, between the S-Graph and A-Graph at room and wall-surface levels, accounting for potential deviations. In Stage-B, the Local Graph Matching extends the previously matched elements in the diS-Graph with newly detected elements following an incremental approach.
@@
Graph Coupling (Section. V). It is performed for the candidates matched by graph matching. In Stage-A, graph coupling registers the origins of two graphs along with the semantic coupling (wall-surfaces and rooms), and explicit mapping of the deviation factors. Stage-B only performs the semantic coupling that incorporates the newly observed entities with explicit deviation factors.
@@
Graph Optimization (Section. VI). It optimizes the coupled graph. Stage-A involves two steps: alternating optimization and joint optimization. Stage-B only does joint optimization. Alternating optimization estimates two types of variables: 1. The ransformation between the origins of the S-Graph and the A-Graph yielding global localization, 2. The possible deviations between the matched graph entities. In joint optimization the coupled graphs are jointly optimized given the initial global transformation and deviation estimates.
@@ IV Graph Matching
Our graph matching extends the method presented in [8]. In [8], a top-down potential candidate search between the A-Graph and S-Graph is performed by leveraging their hierarchical structures. To assess the overall consistency of each generated candidate, two verification steps are applied iteratively. First, the consistency of the node type and graph structure is verified. Second, geometric consistency (i.e. norm) is maintained over a certain consistency threshold [26].
@@ IV-A Deviation-based Candidates Search
Non-Deviated Elements Search To handle potential deviations, our method relaxes the matching criteria by decreasing the consistency thresholds at each level. First, we apply relaxed consistency thresholds for the generation of room-to-room and room-to-wall-surface match candidates, to account for the induced room-center inconsistencies. Then, to exclude deviated wall-surfaces, we increase the threshold for wall-surface-to-wall-surface candidates.
Deviated Elements Detection To identify the deviated wall-surfaces which were not matched in the first stage but are connected to already matched rooms, we decrease the consistency threshold at wall-surface-to-wall-surface level.
@@
Finally, the geometric consistency score provides a quantification of the probability of the deviation for each room and wall-surface, which is further used in the Graph Coupling step (Section V).
@@ IV-B Incremental Matching
To enhance the efficiency of the Graph Matching algorithm in [8], we propose an incremental approach with two stages (associated with stages A and B of the systems architecture of Fig. 2), namely Global Graph Matching (Fig. 4 left) and Local Graph Matching (Fig. 4 right), each executing the two previously described deviation-based candidate search stages. Until a first unique match has been found, the Global Graph Matching is executed for every new observation in the S-Graph. Afterward, the Local Graph Matching is executed every time the diS-Graph is updated with newly observed rooms and wall surfaces. Here, already-matched elements are excluded from candidate generation, and each assessment of intra-level consistency considers the previously matched elements at the corresponding level.
V Graph Coupling
Origin Coupling. We couple the origins of two graphs by introducing a transformation factor . The cost function is defined as:
@@ V Graph Coupling
Semantic Coupling. We couple the wall-surfaces and rooms of the A-Graph and the S-Graph to estimate deviations between the two graphs by introducing deviation factors as follows:
Room Coupling: To estimate the deviation in the pose of a matched room between the two graphs, we define a deviation factor between the two rooms as , where the cost function is defined as:
@@
Wall-Surface Coupling: After coupling the rooms of the A-Graph and the S-Graph, we then couple the wall-surfaces of the coupled rooms. We define the deviation factor between two wall surfaces as . The cost function to estimate the deviation value is defined as: